Help please!
I tried to solve these but I'm not sure if its correct.

4x – 3y = 1
X = 1 – y
4/7 and y = 3/7


8 x – 6y = 0
X + y = 134
x = 39/28 and y = 13/7

I'm also having trouble solving these. Can you please help? (you can use any convenient method)
-3 - 7z = 27,
-6x + y - 5z = `45
4x + 4y - 6z =0
x + 4y + z = 11
3x + 4y = 7
5x + 2y + 2z =19
-5X +5y + z = 8
9x + 8y + 9z = 116
x + 2y + 9z = 86
Thank you in advance!

Answer by edjones(8007)

4x – 3y = 1
x = 1 – y
.
4(1-y)-3y=1
4-4y-3y=1
-7y=-3
y=3/7
x=1- 3/7
x=4/7
.
Ed